In the equation, 3x-12y=6, if y= -2, what is x?
Thanks!

Answer by stanbon(75887)

In the equation, 3x-12y=6, if y= -2, what is x?
-----
3x - 12*-2 = 6
3x + 24 = 6
3x = -18
x = -6
=========
Cheers,
Stan H.
============